MATH 253  Section 501 – 503  P. Yasskin  Fall 1996

ENGINEERING MATHEMATICS III INFORMATION SHEET

INSTRUCTOR:  Dr. Philip B. Yasskin
OFFICE:  Blocker 620 I
OFFICE HOURS: MW 10:15-11:15 in BLOC 620 I

Th 10:25-11:00 in BLOC 126 or by appointment

OFFICE PHONE: 845-3734
MESSAGES:  845-3261 LEAVE YOUR PHONE NUMBER!
E-MAIL:  yasskin@math.tamu.edu
LECTURE:  MWF 9:10 – 10:00 HELD 113
LABS:  501 TTh 8:00-8:50 Tue in BLOC 164 Th (Maple) in BLOC 127
502 TTh 9:35-10:25 Tue in ENPH 216 Th (Maple) in BLOC 126
503 TTh 3:55-4:45 Tue in ENPH 207 Th (Maple) in BLOC 126
REQUIRED TEXTS:  Stewart – Calculus, Third Edition

Belmonte & Yasskin – Vector Calculus with Maple V, Draft

RECOMMENDED TEXT: Boggess, Barrow, Belmonte, Massoud, Morgan,

Rahe, Stecher, Yasskin – CalcLabs with Maple V

RECOMMENDED SOFTWARE:  Maple V, Release 3, Student Edition

with Maple Flight Manual

GRADING: COVERS: POINTS: EXAM DATES:
EXAM 1  Ch. 11  100  Fri 9/27/96
EXAM 2  Ch. 12  100  Fri 10/18/96
EXAM 3  Ch. 13  100  Fri 11/15/96
FINAL  Ch. 11 – 14 200  Mon 12/16/96, 8 – 10 AM, HELD 113
Maple Project 1  50  To be announced
Maple Project 2  50  To be announced
Maple Quiz  50  To be announced
LAB/HW/QUIZ  100
TOTAL  750

I may curve any grade or the total and will then compute the course grade from the following table:

A = 675 – 750 points  D = 450 – 524 points
B = 600 – 674 points  F = 0 – 449 points
C = 525 – 599 points

DESCRIPTION: This is a third course in calculus for engineering majors and covers chapters 11 through 14 of Stewart. In addition we will use a symbolic computer language called MAPLE.

CATALOGUE DESCRIPTION: 253. Engineering Mathematics III. (3-2). Credit 4. I, II, S Vector calculus; calculus of functions of several variables, partial derivatives, directional derivatives, gradient, multiple integration, Green’s and Stokes’ theorems. Prerequisite: MATH 161 or equivalent (e.g. MATH 152).

OTHER POLICIES

1. HOMEWORK will be collected at the assigned class period. Some problems may

be graded but you will not know which ones. Late homework will NOT be accepted.

QUIZZES will be given in lecture or lab and may not be announced. There will be NO make-ups for quizzes.

LAB REPORTS will be collected at the lab period one week after the lab. Late lab reports will NOT be accepted. On lab reports, students will work in pairs. Each pair will turn in one lab report and receive one grade.

Homeworks, quizzes and lab reports will each count equally. The lowest two homework or quiz or lab report grades will be dropped. The remaining grades will be averaged and then rescaled to 100 points.

2. MAKE-UPS for MAJOR EXAMS will be given only in case of an absence authorized under University Regulations. You will need a note from your doctor or your academic advisor. If you know in advance that you will miss an exam, please contact me so that you can take the make-up in advance. If you have trouble reaching me, leave a message with the Math department secretaries (845-3261) and be sure to leave your phone number.

3. ATTENDANCE is REQUIRED. Attendance will be taken in lecture. If you sign the roll sheet, you are expected to remain in the classroom for the entire 50 minutes. More than 2 absences may have a detrimental effect on your grade especially in borderline cases.

4. You may be asked to provide BLUE BOOKS and multiple choice SCANTRON forms. You must have your ID with you at all exams. You MAY use a simple CALCULATOR during exams but NO PROGRAMMABLE, GRAPHICS or ALGEBRAIC CALCU-LATORS and NO LAPTOP COMPUTERS. You will use the COMPUTER during the Maple quiz.

5. All material handed out or written on the board or spoken in class is copyrighted by the instructor.